OpenStudy (anonymous):

In right ∆ABC, angle B = 90°, angle C = 40°, and angle BC = 10. What are the other two side lengths of the triangle? A) AC = 13.1, AB = 8.4 B) AC = 18.7, AB = 15.8 C) AC = 14.3, AB = 10.2 D) AC = 15.6, AB = 11.9

OpenStudy (anonymous):

Could really use the help, don't have to give me the answer outright just need to know how to solve this problem, can someone please guide me in the right direction? thank you.

Nnesha (nnesha):

 \[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\]